A parked police car takes off after a suspect and reaches 80.2 km/h E in 8.16 seconds. What is the average acceleration (m/s)^2 of the car?
Answer #1

Final speed of the car, v = 80.2 km/h

Covert the units of speed in m/s.

v = (80.2 x 1000) / (60 x 60) = 22.28 m/s

Average acceleration = change in velocity / change in time

= (22.28 - 0) / 8.16 = 22.28 / 8.16 = 2.73 m/s^2 (Answer)